# Recidivism and Peer Influence with LLM Text Embeddings in Low Security Correctional Facilities

### v1

@@ 3.1 LLM embedding of text and Zero-Shot classification
We deploy Pre-trained LLMs on the corpus of written messages to derive two representations: (1) context-sensitive embedding vectors and (2) probability vectors for meaningful psychological categories using Zero-Shot classification. We use the bert-base-uncased model (Devlin et al., 2019) for (1), and the bart-large model (Lewis et al., 2019) for (2). Before detailing the model structure, we briefly review the development of text embedding methods and discuss our rationale for adopting transformer-based AI models, such as BERT and BART.
A text embedding is a high-dimensional numerical vector representation of text. If two words convey similar meanings, they are represented closely in the vector space. However, earlier word embedding models, such as Word2Vec (Mikolov et al., 2013), did not consider the context for words, limiting their ability to handle words with multiple context-specific meanings. Transformer-based models, like BERT and BART, overcome these limitations using self-attention mechanisms (Vaswani et al., 2017), where the embedding for each word is dynamically influenced by the embedding of the entire input sequence, providing potentially different embeddings for each word based on context. The generative pre-training (GPT) framework Radford et al. (2018) uses semi-supervised learning in the transformer architecture. The model is pre-trained unsupervised on a large data corpus to learn embeddings or representations, followed by supervised fine-tuning for specific tasks. The language representation model BERT in Devlin et al. (2019) further improved the capabilities of pre-trained language models by introducing a bi-directional transformer, which can use both left and right context for a word as opposed to unidirectional processing (typically from left to right) in OpenAI GPT-1 of Radford et al. (2018).
@@
A sequence of transformer blocks processes $\mathbf{E}$, where the output of each block is fed as an input into the next block. Let the output of the $l$-th block be $\mathbf{E}^{(l)}\in\mathbb{R}^{K\times H}$, with the initial input being $\mathbf{E}^{(0)}=\mathbf{E}$. We next describe the processes involved in each of these transformer blocks and the multi-head self-attention mechanism.
Multi-Head Self-Attention and transformer blocks
The query vector represents what the $k$-th token seeks, and the key vector indicates what each token offers others. The attention score between the $k$-th query and all $K$ keys is given by $\mathbf{s}_{k}^{(l,a)}=\left[s_{k1}^{(l,a)},s_{k2}^{(l,a)},\ldots,s_{kK}^{(l,a)}\right]^{\top}\in\mathbb{R}^{K}$. Each element in this vector is the scaled dot-product between the corresponding query and key vectors, $s_{kj}^{(l,a)}=\frac{\mathbf{q}_{k}^{(l,a)\top}\mathbf{k}_{j}^{(l,a)}}{\sqrt{d}}.$ The attention weights are then obtained by normalizing $s_{kj}^{(l,a)}$ using the softmax function as $m_{kj}^{(l,a)}=\frac{\exp(s_{kj}^{(l,a)})}{\sum_{q=1}^{K}\exp(s_{kq}^{(l,a)})}.$ Let $\mathbf{M}^{(l,a)}\in\mathbb{R}^{K\times K}$ be the attention weights matrix for all tokens in $\mathbf{z}$. The output matrix is given by $\mathbf{H}^{(l,a)}=\mathbf{M}^{(l,a)}\mathbf{V}^{(l,a)}\in\mathbb{R}^{K\times d}$. Next, the outputs from $A$ heads are combined as: $\mathbf{H}^{(l)}=\left[\mathbf{H}^{(l,1)},\mathbf{H}^{(l,2)},\ldots,\mathbf{H}^{(l,A)}\right]\in\mathbb{R}^{K\times H}$. The column-concatenated matrix $\mathbf{H}^{(l)}$ is then processed within the encoder block to generate the block’s final output $\mathbf{E}^{(l)}$ (Figure B5 in Appendix).
Text Embedding using BERT
To numerically encode the semantic content of each message, we use the output from the 11th encoder block (i.e., the penultimate block) of bert-base-uncased model to construct an “embedding profile” for each participant — a high-dimensional vector summarizing their language use across multiple messages during their stay in TC. We choose the penultimate layer as that layer is widely thought to contain the unsupervised representation of data in the representation learning framework (Bengio et al., 2013), while the last layer is trained or fine-tuned to be task-specific. One can, in principle, also choose an embedding representation from an earlier layer, but not the last layer. Let $\mathbf{f}^{\mathrm{token}}_{c,k}\in\mathbb{R}^{H}$ denote the embedding of the $k$-th token in the $c$-th message, extracted from the 11th encoder. We obtain the message-level embedding by taking the average of these vectors over all tokens in the message as $\mathbf{f}^{\mathrm{msg}}_{c}=\frac{1}{N_{c}}\sum_{k=1}^{N_{c}}\mathbf{f}^{\mathrm{token}}_{c,k}$, where $N_{c}$ is the number of tokens in message $c$. We then compute the “sender embedding profile” for a sender $s$ by taking an average of the the message-level embeddings, $\mathbf{f}^{\mathrm{sender}}_{s}=\frac{1}{N_{s}}\sum_{c\in\mathcal{C}_{s}}\mathbf{f}_{c}^{\mathrm{msg}}$, where $\mathcal{C}_{s}$ denotes the set of messages sent by sender $s$, and $N_{s}=|\mathcal{C}_{s}|$ is the number of such messages. The embedding profile for each receiver can be obtained analogously.
@@ 3.2 Zero-Shot Classification
To interpret the embeddings, we relate them to interpretable psychological constructs, using the “Zero-Shot” classification framework (Yin et al., 2019). In Zero-Shot classification, a user provides an arbitrary number of labels, and the LLM assigns given texts to the most appropriate label. This enables us to assign written affirmations and corrections to psychologically relevant categories without requiring additional training.
@@ 3.3 Prediction models
We fit a penalized regression lasso method (Tibshirani, 1996) to predict recidivism using the resident-level covariates along with the individual embedding profiles. Suppose $Y_{i}$ denotes the recidivism status for individual $i$, which takes the value $1$ if the individual recidivates and 0 if the person does not recidivate. Therefore, we can model $Y_{i}$ with a Bernoulli distribution with probability of success being $p_{i}$ and then further model $p_{i}$ using the covariates and embedding profiles. However, since the embedding profiles are high-dimensional, we use a penalized logistic regression method (logistic-lasso). Suppose $\bm{X}$ is the $N\times p$ matrix of observed covariates and $\bm{T}$ is the $N\times H$ matrix containing the dimensions of text embeddings obtained from the LLM model as columns. Then the logistic regression model is $\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\bm{X}_{i}\beta_{1}+\bm{T}_{i}\beta_{2}$, where $\beta_{1}$ and $\beta_{2}$ are $p$ and $H$ dimensional parameter vectors respectively. We optimize for $\beta_{1},\beta_{2}$ by maximizing the penalized likelihood function with $\ell_{1}$ penalty as follows,
@@
Beyond predicting recidivism, we also want to interpret the link between individual behavioral and emotional profiles as manifested by the messages with recidivism. In order to do so we perform a Zero-Shot classification as described before to classify the messages into the four groups. Let $\bm{Q}$ denote the $N\times 4$ matrix containing the average of probabilities assigned to the messages sent (respectively received) by each individual. Since the predictors are now low-dimensional (just four dimensions in addition to the covariates), we will not need to penalize the coefficients. However, we note that some of the covariates are now “compositional”, since $\sum_{j}Q_{ij}=1$ for each resident $i$. Therefore, following standard techniques of dealing with compositional predictors, we designate one covariate as baseline and take log ratios (ALR) of other covariates with respect to the baseline covariate (Aitchison, 1982). This is important for interpreting the model coefficients, but not for prediction accuracy. We note that we could just continue with the original predictors and drop the intercept, and have a model with comparable prediction accuracy.
3.4 Multivariate Peer Effects Model with Endogenous Network Formation
@@
where $\mathbf{Y}\in\mathbb{R}^{N\times m}$ is the matrix of outcome variables, $\mathbf{G}=(g_{ij})\in\mathbb{R}^{N\times N}$ where $g_{ij}:=\frac{a_{ij}}{\sum_{j\neq i}a_{ij}}$, is the row-normalized counterpart of the network adjacency matrix $\mathbf{A}$, and $\mathbf{X}\in\mathbb{R}^{N\times p}$ is the matrix of observed covariates. We write $\bm{Y}_{i}\in\mathbb{R}^{m}$ and $\bm{Y}_{.,j}\in\mathbb{R}^{N}$ to denote the $i$-th row and $j$-column vectors of an $N\times m$ matrix $\mathbf{Y}$, respectively. The parameters have the following interpretation. The matrix $\mathbf{D}\in\mathbb{R}^{m\times m}$ is the peer effect parameter matrix. The diagonal of $\mathbf{D}$ are the direct peer spillover effects on the same dimensions of $\mathbf{Y}$, and the off-diagonal elements are the indirect peer spillover effects of a dimension on another dimension. The matrices $\mathbf{B}_{1},\mathbf{B}_{2}\in\mathbb{R}^{p\times m}$ are coefficient matrices for $\mathbf{X}$ and $\mathbf{G}\mathbf{X}$, respectively. However, since the network is endogenous, we cannot identify and consistently estimate $\mathbf{D}$ from this model using Bramoullé et al. (2009)’s IV 2LSLS framework. For concreteness, we assume for each individual there is a vector $\mathbf{U}_{i}$ of latent homophily variables that is correlated with both the network adjacency matrix $\mathbf{A}$ and the error matrix $\mathbf{E}$. We assume the error matrix $\mathbf{E}\in\mathbb{R}^{N\times m}$ is such that for each row $\mathbb{E}[\mathbf{E}_{i}|\mathbf{U}_{i}]=\bm{h}^{E}(\bm{U}_{i})$ for some unknown function $\bm{h}^{E}$, and $\mathbb{V}[\mathbf{E}_{i}|\mathbf{U}_{i}]=\mathbf{V}$. The matrix $\mathbf{V}$ allows dependence across outcome dimensions. The vectors $\mathbf{E}_{i},\mathbf{X}_{i},\mathbf{U}_{i}$ are assumed to be i.i.d. across individuals $i$. We let $A_{ij}=f(\mathbf{U}_{i},\mathbf{U}_{j},\xi_{ij})$, where $\xi_{ij}$s over all $(i,j)$ are assumed to be i.i.d and independent of $\mathbf{X,E,U}$. We postpone the discussion on models for network formation until the next section. We further assume that $\mathbb{E}[\bm{E}_{i}|\bm{X}_{i},\bm{U}_{i}]=\mathbb{E}[\bm{E}_{i}|\bm{U}_{i}]$, i.e., conditional on $\bm{U}_{i}$, the covariates $\bm{X}_{i}$ and error term $\bm{E}_{i}$ are uncorrelated.
@@
This model is similar to the multivariate extension of the spatial autoregressive (MSAR) model proposed in (Zhu et al., 2020). However, the model in Zhu et al. (2020) can be used for peer effect estimation only under the assumption of exogenous network formation, i.e., $\mathbf{G}$ is uncorrelated with $\mathbf{E}$. Further, model (1) can also be thought of as a multivariate extension of the endogenous and simultaneous peer effect model in Johnsson and Moon (2021), both in terms of multidimensional outcome variables ($\mathbf{Y}$) and latent variables ($\mathbf{U}$).
@@
To identify the parameter matrices, we take an instrumental variable approach similar to Johnsson and Moon (2021), but extend the methodology in several directions as we detail below. We define the combined regressor and instrument matrices as:
@@
This estimator is a multivariate extension of the estimator proposed in Bramoullé et al. (2009); Kelejian and Prucha (1998) and is a consistent estimator provided the network is exogenous, which is not the case in our setup. Taking expectations on both sides of Equation 1 conditioning on $\bm{U}$ and subtracting this from equation 1, we obtain
@@
The following lemma provides sufficient conditions for the identification of the true parameter matrix $\beta_{0}$.
Lemma 1.
@@
Therefore, given access to the latent positions $\mathbf{U}$ and the three conditional mean functions, the parameter $\boldsymbol{\beta}_{0}$ can be estimated using 2SLS. Note that the rank condition in Lemma 1 can generally be satisfied with $p>m$, i.e., if we have more covariates than dimensions. However, the moment equation cannot be estimated as we do not observe $\mathbf{U}$ and do not know the conditional expectations of $\mathbf{Y,Z,K}$ given $\mathbf{U}$. We next address how to estimate both of these.
@@ 3.5 Latent positions and Non-Parametric estimation with Sieves
As mentioned earlier, we model the network data using a latent variable model $A_{ij}=f(\mathbf{U}_{i},\mathbf{U}_{j},\xi_{ij})$ and posit that the node level latent variables $\mathbf{U}_{i}$ are involved in both the network formation model and are correlated with the error term of the outcome model. These latent variables represent unobserved characteristics of individuals, which may be responsible for tie formation (latent homophily), and an unknown function of these latent variables is part of the error term in the outcome model. Therefore, these latent variables can be estimated from the observed network. This is a core assumption made in recent methodological advances in peer effect estimation (McFowland III and Shalizi, 2023; Johnsson and Moon, 2021; Nath et al., 2025; Goldsmith-Pinkham and Imbens, 2013). However, while McFowland III and Shalizi (2023); Nath et al. (2025) focused on the longitudinal peer effect model, we consider the simultaneous peer effect model. Further, McFowland III and Shalizi (2023); Nath et al. (2025) assumed the latent variables enter the model for $\mathbf{Y}$ linearly as $\mathbf{U\beta}$. In contrast, we assume the conditional expectations of $\mathbf{Y,Z,K}$ to be unknown functions $\mathbf{h(U)}$ of the latent variables. In this aspect, our approach resembles Johnsson and Moon (2021), however, we consider more general latent variable models for network data with multi-dimensional latent variables which are more realistic for modeling individual characteristics. Further, the theoretical results in Johnsson and Moon (2021) requires the network to be “dense” in the sense that every individual is expected to be connected to $O(N)$ individuals and consequently, total number of edges in the network is expected to be $O(N^{2})$. This is quite unrealistic as real-world networks tend to be sparse with individuals connected to only a few other people, even when the network size is huge (e.g., in online social networks, even if there are millions of people present in the network, an individual member only has a few hundred or thousands of connections). In contrast, our results here hold for sparse networks.
@@
To estimate each component function $h_{j}(\bm{u})$, we use tensor-product basis functions Zhang and Simon (2023) given by $\phi_{k}(\bm{u})=\prod_{\ell=1}^{d}\psi_{k_{\ell}}^{(\ell)}(u_{\ell})$, where each $\psi_{k_{\ell}}^{(\ell)}$ is a univariate basis function (polynomial or cosine) on the $\ell$-th coordinate of $\mathbf{u}$. With these basis functions, we approximate each component function $h_{j}(\bm{u})$, by a linear combination of basis functions:
@@
Since latent positions $\bm{U}$ are unknown, we use a two-step procedure. For estimating $\bm{U}$, we model the network using latent variable models such as the RDPG model (Athreya et al., 2017; Rubin-Delanchy et al., 2022; Xie and Xu, 2023) and the additive and multiplicative effects latent space model (Ma et al., 2020; Hoff, 2021; Hoff et al., 2002; Li et al., 2023). Then our two-stage procedure is as follows. (i) we first construct an estimator $\widehat{\bm{U}}:=(\widehat{\bm{u}}_{1},\dots,\widehat{\bm{u}}_{N})^{\top}$ for the latent traits using spectral embedding for RDPG models or maximum likelihood estimation for additive and multiplicative effects latent space models, and (ii) we evaluate the basis functions at $\widehat{\bm{U}}$ and apply the same projection strategy. With these estimated latent vectors, we define the estimated design matrix as $\widehat{\mathbf{\Phi}}_{N}:=\mathbf{\Phi}_{N}(\widehat{\bm{U}})=\left[\bm{\phi}^{L_{N}}(\widehat{\bm{u}}_{1}),\ldots,\bm{\phi}^{L_{N}}(\widehat{\bm{u}}_{N})\right]\in\mathbb{R}^{N\times L_{N}}.$ The final estimated conditional mean functions are $\hat{\bm{h}}^{W}(\widehat{\bm{U}}):=\mathbf{P}_{\widehat{\mathbf{\Phi}}_{N}}\bm{W}$ where $\mathbf{P}_{\widehat{\mathbf{\Phi}}_{N}}:=\widehat{\mathbf{\Phi}}_{N}(\widehat{\mathbf{\Phi}}_{N}^{\top}\widehat{\mathbf{\Phi}}_{N})^{-1}\widehat{\mathbf{\Phi}}_{N}^{\top}$.
Denote $\bm{M}_{\hat{\bm{\Phi}}_{N}}=\bm{I}_{N}-\bm{P}_{\hat{\mathbf{\Phi}}_{N}}$. Then, our Two-Stage Least Squares (2SLS) estimator is defined as:

### v2

@@ 4.1 LLM embedding of text and Zero-Shot classification
We deploy Pre-trained LLMs on the corpus of written messages to derive two representations: (1) context-sensitive embedding vectors and (2) probability vectors for meaningful psychological categories using Zero-Shot classification. We use the bert-base-uncased model (Devlin et al.,, 2019) for (1), and the bart-large model (Lewis et al.,, 2019) for (2). Before detailing the model structure, we briefly review the development of text embedding methods and discuss our rationale for adopting transformer-based AI models, such as BERT and BART.
A text embedding is a high-dimensional numerical vector representation of text. If two words convey similar meanings, they are represented closely in the vector space. However, earlier word embedding models, such as Word2Vec (Mikolov et al.,, 2013), did not consider the context for words, limiting their ability to handle words with multiple context-specific meanings. Transformer-based models, like BERT and BART, overcome these limitations using self-attention mechanisms (Vaswani et al.,, 2017), where the embedding for each word is dynamically influenced by the embedding of the entire input sequence, providing potentially different embeddings for each word based on context. The generative pre-training (GPT) framework Radford et al., (2018) uses semi-supervised learning in the transformer architecture. The model is pre-trained unsupervised on a large data corpus to learn embeddings or representations, followed by supervised fine-tuning for specific tasks. The language representation model BERT in Devlin et al., (2019) further improved the capabilities of pre-trained language models by introducing a bi-directional transformer, which can use both left and right context for a word as opposed to unidirectional processing (typically from left to right) in OpenAI GPT-1 of Radford et al., (2018).
@@
A sequence of transformer blocks processes $\mathbf{E}$, where the output of each block is fed as an input into the next block. Let the output of the $l$-th block be $\mathbf{E}^{(l)}\in\mathbb{R}^{K\times H}$, with the initial input being $\mathbf{E}^{(0)}=\mathbf{E}$. We review the processes involved in each of these transformer blocks and the multi-head self-attention mechanism in section A4 of the Appendix.
To numerically encode the semantic content of each message, we use the output from the 11th encoder block (i.e., the penultimate block) of bert-base-uncased model to construct an “embedding profile” for each participant — a high-dimensional vector summarizing their language use across multiple messages during their stay in TC. We choose the penultimate layer as that layer is widely thought to contain the unsupervised representation of data in the representation learning framework (Bengio et al.,, 2013), while the last layer is trained or fine-tuned to be task-specific. One can, in principle, also choose an embedding representation from an earlier layer. Let $\mathbf{f}^{\mathrm{token}}_{c,k}\in\mathbb{R}^{H}$ denote the embedding of the $k$-th token in the $c$-th message, extracted from the 11th encoder. We obtain the message-level embedding by taking the average of these vectors over all tokens in the message as $\mathbf{f}^{\mathrm{msg}}_{c}=\frac{1}{N_{c}}\sum_{k=1}^{N_{c}}\mathbf{f}^{\mathrm{token}}_{c,k}$, where $N_{c}$ is the number of tokens in message $c$. We then compute the “sender embedding profile” for a sender $s$ by taking an average of the the message-level embeddings, $\mathbf{f}^{\mathrm{sender}}_{s}=\frac{1}{N_{s}}\sum_{c\in\mathcal{C}_{s}}\mathbf{f}_{c}^{\mathrm{msg}}$, where $\mathcal{C}_{s}$ denotes the set of messages sent by sender $s$, and $N_{s}=|\mathcal{C}_{s}|$ is the number of such messages. The embedding profile for each receiver can be obtained analogously.
@@ 4.2 Zero-Shot Classification
To interpret the embeddings, we relate them to interpretable psychological constructs, using the “Zero-Shot” classification framework (Yin et al.,, 2019). In Zero-Shot classification, a user provides an arbitrary number of labels, and the LLM assigns given texts to the most appropriate label. This enables us to assign written affirmations and corrections to psychologically relevant categories without requiring additional training.
@@ 4.3 Prediction models
We fit a penalized regression lasso method (Tibshirani,, 1996) to predict recidivism using the resident-level covariates along with the individual embedding profiles. Suppose $Y_{i}$ denotes the recidivism status for individual $i$, which takes the value $1$ if the individual recidivates and 0 if the person does not recidivate. Therefore, we can model $Y_{i}$ with a Bernoulli distribution with probability of success being $p_{i}$ and then further model $p_{i}$ using the covariates and embedding profiles. However, since the embedding profiles are high-dimensional, we use a penalized logistic regression method (logistic-lasso). Suppose $\bm{X}$ is the $N\times p$ matrix of observed covariates and $\bm{T}$ is the $N\times H$ matrix containing the dimensions of text embeddings obtained from the LLM model as columns. Then the logistic regression model is $\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\bm{X}_{i}\beta_{1}+\bm{T}_{i}\beta_{2}$, where $\beta_{1}$ and $\beta_{2}$ are $p$ and $H$ dimensional parameter vectors respectively. We optimize for $\beta_{1},\beta_{2}$ by maximizing the penalized likelihood function with $\ell_{1}$ penalty as follows,
@@
Beyond predicting recidivism, we also want to interpret the link between individual behavioral and emotional profiles as manifested by the messages with recidivism. In order to do so we perform a Zero-Shot classification as described before to classify the messages into the four groups. Let $\bm{Q}$ denote the $N\times 4$ matrix containing the average of probabilities assigned to the messages sent (respectively received) by each individual. Since the predictors are now low-dimensional (just four dimensions in addition to the covariates), we will not need to penalize the coefficients. However, we note that some of the covariates are now “compositional”, since $\sum_{j}Q_{ij}=1$ for each resident $i$. Therefore, following standard techniques of dealing with compositional predictors, we designate one covariate as baseline and take log ratios (ALR) of other covariates with respect to the baseline covariate (Aitchison,, 1982). This is important for interpreting the model coefficients, but not for prediction accuracy.  [^5] We note that we could just continue with the original predictors and drop the intercept, and have a model with comparable prediction accuracy.
4.4 Multivariate Peer Effects Model with Endogenous Network
@@
where $\mathbf{Y}\in\mathbb{R}^{N\times m}$ is the matrix of outcome variables, $\mathbf{G}=(g_{ij})\in\mathbb{R}^{N\times N}$ where $g_{ij}:=\frac{a_{ij}}{\sum_{j\neq i}a_{ij}}$, is the row-normalized counterpart of the network adjacency matrix $\mathbf{A}$, and $\mathbf{X}\in\mathbb{R}^{N\times p}$ is the matrix of observed covariates. We write $\bm{Y}_{i}\in\mathbb{R}^{m}$ and $\bm{Y}_{.,j}\in\mathbb{R}^{N}$ to denote the $i$-th row and $j$-column vectors of an $N\times m$ matrix $\mathbf{Y}$, respectively. Matrix $\mathbf{D}\in\mathbb{R}^{m\times m}$ is the peer effect parameter matrix. The diagonal of $\mathbf{D}$ are the direct peer spillover effects on the same dimensions of $\mathbf{Y}$, and the off-diagonal elements are the indirect peer spillover effects of a dimension on another dimension. The matrices $\mathbf{B}_{1},\mathbf{B}_{2}\in\mathbb{R}^{p\times m}$ are coefficient matrices for $\mathbf{X}$ and $\mathbf{G}\mathbf{X}$, respectively.
@@
However, since the network is endogenous, we cannot identify and consistently estimate $\mathbf{D}$ from this model using Bramoullé et al., (2009)’s IV 2LSLS framework. For concreteness, we assume for each individual there is a vector $\mathbf{U}_{i}$ of latent homophily variables that is correlated with both the network adjacency matrix $\mathbf{A}$ and the error matrix $\mathbf{E}$. We assume the error matrix $\mathbf{E}\in\mathbb{R}^{N\times m}$ is such that for each row $\mathbb{E}[\mathbf{E}_{i}|\mathbf{U}_{i}]=\bm{h}^{E}(\bm{U}_{i})$ for some unknown function $\bm{h}^{E}$, and $\mathbb{V}[\mathbf{E}_{i}|\mathbf{U}_{i}]=\mathbf{V}$. The matrix $\mathbf{V}$ allows dependence across outcome dimensions. The vectors $\mathbf{E}_{i},\mathbf{X}_{i},\mathbf{U}_{i}$ are assumed to be i.i.d. across individuals $i$. We let $A_{ij}=f(\mathbf{U}_{i},\mathbf{U}_{j},\xi_{ij})$, where $\xi_{ij}$s over all $(i,j)$ are assumed to be i.i.d and independent of $\mathbf{X,E,U}$. We postpone the discussion on the function $f$ and the models for network formation until the next section. We further assume that $\mathbb{E}[\bm{E}_{i}|\bm{X}_{i},\bm{U}_{i}]=\mathbb{E}[\bm{E}_{i}|\bm{U}_{i}]$, i.e., conditional on $\bm{U}_{i}$, the covariates $\bm{X}_{i}$ and error term $\bm{E}_{i}$ are uncorrelated.
@@
This model is similar to the multivariate extension of the spatial autoregressive (MSAR) model proposed in (Zhu et al.,, 2020). However, the model in Zhu et al., (2020) can be used for peer effect estimation only under the assumption of exogenous network formation ($\mathbf{G}$ is uncorrelated with $\mathbf{E}$) and the proposed estimator is the maximum likelihood estimator and not IV2SLS. Further, model (1) can also be thought of as extending the endogenous and simultaneous peer effect model in Johnsson and Moon, (2021), both in terms of multidimensional outcome and latent variables.
@@
To identify the parameter matrices, we take an instrumental variable approach similar to Johnsson and Moon, (2021), but extend the methodology in several directions as we detail below. We define the combined regressor and instrument matrices as:
@@
This estimator is a multivariate extension of the estimator proposed in Bramoullé et al., (2009); Kelejian and Prucha, (1998) and is a consistent estimator provided the network is exogenous, which is not the case in our setup. Taking expectations on both sides of Equation 1 conditioning on $\bm{U}$ and subtracting this from equation 1, we obtain
@@
The following proposition provides sufficient conditions for the identification of the true parameter matrix $\beta_{0}$.
Proposition 1.
@@
Therefore, given access to the latent positions $\mathbf{U}$ and the three conditional mean functions, the parameter $\boldsymbol{\beta}_{0}$ can be estimated using 2SLS. Note that the rank condition in Proposition 1 can generally be satisfied with $p>m$, i.e., if we have more covariates than dimensions. However, the moment equation still cannot be estimated as we do not observe $\mathbf{U}$ and do not know the conditional expectations of $\mathbf{Y,Z,K}$ given $\mathbf{U}$. We next address how to estimate both of these components.
@@ 4.5 Latent positions and Non-Parametric estimation with Sieves
As mentioned earlier, we model the network data using a latent variable model $A_{ij}=f(\mathbf{U}_{i},\mathbf{U}_{j},\xi_{ij})$ and posit that the node level latent variables $\mathbf{U}_{i}$ are involved in both the network formation model and are correlated with the error term of the outcome model. These latent variables represent unobserved characteristics of individuals, which may be responsible for tie formation (latent homophily), and an unknown function of these latent variables is part of the error term in the outcome model. Therefore, these latent variables can be estimated from the observed network. This is a core assumption made in recent methodological advances in peer effect estimation (McFowland III and Shalizi,, 2023; Johnsson and Moon,, 2021; Nath et al.,, 2025; Goldsmith-Pinkham and Imbens,, 2013). However, while McFowland III and Shalizi, (2023); Nath et al., (2025) focused on the longitudinal peer effect model, we consider the simultaneous peer effect model. Further, McFowland III and Shalizi, (2023); Nath et al., (2025) assumed the latent variables enter the model for $\mathbf{Y}$ linearly as $\mathbf{U\beta}$. In contrast, we assume the conditional expectations of $\mathbf{Y,Z,K}$ to be unknown functions $\mathbf{h(U)}$ of the latent variables. In this aspect, our approach resembles Johnsson and Moon, (2021); however, we consider more general latent variable models for network data with multi-dimensional latent variables, which are more realistic for modeling individual characteristics. Further, the theoretical results in Johnsson and Moon, (2021) require the network to be “dense” in the sense that every individual is expected to be connected to $O(N)$ individuals and consequently, the total number of edges in the network is expected to be $O(N^{2})$. This is quite unrealistic as real-world networks tend to be sparse with individuals connected to only a few other people, even when the network size is huge (e.g., in online social networks, even if there are millions of people present in the network, an individual member only has a few hundred or thousands of connections). In contrast, our results here hold for sparse networks.
@@
To estimate each component function $h_{j}(\bm{u})$, we use tensor-product basis functions Zhang and Simon, (2023) given by $\phi_{k}(\bm{u})=\prod_{\ell=1}^{d}\psi_{k_{\ell}}^{(\ell)}(u_{\ell})$, where each $\psi_{k_{\ell}}^{(\ell)}$ is a univariate basis function (polynomial or cosine) on the $\ell$-th coordinate of $\mathbf{u}$. With these basis functions, we approximate each component function $h_{j}(\bm{u})$, by a linear combination of basis functions:
@@
Since latent positions $\bm{U}$ are unknown, we use a two-step procedure. For estimating $\bm{U}$, we model the network using latent variable models such as the RDPG model (Athreya et al.,, 2018; Rubin-Delanchy et al.,, 2022; Xie and Xu,, 2023) and the additive and multiplicative effects latent space model (Ma et al.,, 2020; Hoff,, 2021; Hoff et al.,, 2002; Li et al.,, 2023). Then our two-stage procedure is as follows. (i) we first construct an estimator $\widehat{\bm{U}}:=(\widehat{\bm{u}}_{1},\dots,\widehat{\bm{u}}_{N})^{\top}$ for the latent traits using spectral embedding for RDPG models or maximum likelihood estimation for additive and multiplicative effects latent space models, and (ii) we evaluate the basis functions at $\widehat{\bm{U}}$ and apply the same projection strategy. With these estimated latent vectors, we define the estimated design matrix as $\widehat{\mathbf{\Phi}}_{N}:=\mathbf{\Phi}_{N}(\widehat{\bm{U}})=\left[\bm{\phi}^{L_{N}}(\widehat{\bm{u}}_{1}),\ldots,\bm{\phi}^{L_{N}}(\widehat{\bm{u}}_{N})\right]\in\mathbb{R}^{N\times L_{N}}.$ The final estimated conditional mean functions are $\hat{\bm{h}}^{W}(\widehat{\bm{U}}):=\mathbf{P}_{\widehat{\mathbf{\Phi}}_{N}}\bm{W}$ where $\mathbf{P}_{\widehat{\mathbf{\Phi}}_{N}}:=\widehat{\mathbf{\Phi}}_{N}(\widehat{\mathbf{\Phi}}_{N}^{\top}\widehat{\mathbf{\Phi}}_{N})^{-1}\widehat{\mathbf{\Phi}}_{N}^{\top}$.
Denote $\bm{M}_{\hat{\bm{\Phi}}_{N}}=\bm{I}_{N}-\bm{P}_{\hat{\mathbf{\Phi}}_{N}}$. Then, our Two-Stage Least Squares (2SLS) estimator is: